# The Infrared Absorption Spectrum of Phenylacetylene and its Deuterated Isotopologue in the Mid- to Far-IR

### 24 Jan 2024 v1

@@ Abstract
Anharmonicity strongly influences the absorption and emission spectra of polycyclic aromatic hydrocarbon (PAH) molecules. Here, ion-dip spectroscopy experiments together with detailed anharmonic computations reveal the presence of fundamental, overtone, as well as 2- and 3-quanta combination band transitions in the far- and mid-infrared absorption spectrum of phenylacetylene and its singly deuterated isotopologue. Strong absorption features in the 400-900 cm range originate from CH(D) in-plane and out-of-plane (OOP) wags and bends, as well as bending motions including the CC and CH bonds of the acetylene substituent and the aromatic ring. For phenylacetylene, every absorption feature is assigned either directly or indirectly to a single or multiple vibrational mode(s). The measured spectrum is dense, broad, and structureless in many regions but well characterized by computations. Upon deuteration, large isotopic shifts are observed. At frequencies above 1500 cm for d–phenylacetylene, a one-to-one match is seen when comparing computations and experiment with all features assigned to combination bands and overtones. The CC stretch observed in phenylacetylene is not observed in d–phenylacetylene due to a computed 40-fold drop in intensity. Overall, a careful treatment of anharmonicity that includes 2- and 3-quanta modes is found to be crucial to understand the rich details of the infrared spectrum of phenylacetylene. Based on these observations it can be expected that such an all-inclusive anharmonic treatment will also be key for unraveling the infrared spectra of PAHs in general.
@@ I Introduction
PAHs are a class of molecules composed of fused hexagonal rings of sp-hybridized carbon atoms with hydrogens decorating the periphery, which are known to be ubiquitous in space based on their infrared emission features detected throughout many interstellar regions, i.e., the PAH bands.[1, 2] Recently, experimental and computational studies of the infrared absorption spectra of various PAHs have revealed large anharmonic effects. Previous attempts to account for anharmonicity included the use of harmonic scaling factors.[3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] However, scaling factors do not account for important higher-order transitions such as combination bands and overtones, nor do they provide a good treatment of systems with shallow PESs.[14]
@@
In the recent past dedicated gas-phase experiments have been performed for a number of PAHs, and these indeed reveal that for these compounds anharmonicity is essential to come to a proper interpretation of their infrared absorption spectra. For example, ion-dip molecular beam experiments have been used to record infrared absorption spectra of neutral naphthalene, anthracene, tetracene and pentacene, showing the appearance of many combination bands.[25, 17, 18, 26, 27, 28] Similarly, molecular beam experiments on the larger PAHs coronene, peropyrene, ovalene, and hexa(peri)benzocoronene, have shown the presence of many combination bands and overtones in the 1600–2000 cm and 2950–3150 cm regions, whereas at lower frequencies (100 and 1000 cm) fundamental transitions dominate.[29] Moreover, storage ring experiments on laser-excited cationic cyanonaphthalene have shown rapid cooling via recurrent fluorescence, with computations only agreeing if anharmonicity is considered.[30] Recently, using an infrared table-top laser system to cover the 3 µm spectral region, the same research group performed ion-dip experiments on phenylacetylene and d–phenylacetylene (the singly deuterated isotopologue of phenylacetylene with deuteration taking place at the acetylene hydrogen),[16] representing an ideal model system to assess the impact of anharmonic effects on PAHs, in particular focusing on the possibility that higher-order quanta modes are a significant contributor to their infrared spectra. Remarkably, a major influence of 2- and 3-quanta modes arising from aromatic and acetylenic C-H and C-D stretching bands is observed.[31] Notably, phenylacetylene was recently detected in the dark molecular cloud TMC-1 via its pure rotational transitions,[32] making the detailed understanding of its infrared spectrum a pressing need.
@@ II.1 Experimental
Molecular beams of phenylacetylene and d–phenylacetylene are formed by placing the liquid compounds in a reservoir kept at room temperature that is seeded with a stream of Ar gas injected at a backing pressure of 4 bar. Using a Series 9 pulsed valve from General Valve, the Ar-phenylacetylene mixtures are expended into vacuum, forming a cold and directional molecular beam that is collimated by a 2 mm skimmer. The pulsed valve is operated at 10 Hz. After collimation, the molecular beam enters the ionization region of a reflectron time-of-flight mass spectrometer (R.M. Jordan D-850) equipped with a 40 mm dual microchannel plate (MCP) detector (Jordan Co. C-726) that has a mass resolution of = 2200 at 100 amu.
@@ II.2 Computational
Computation of the optimized geometry; harmonic normal modes; and quadratic, cubic, and quartic normal coordinate force constants (quartic force field; QFF) is performed at the B3LYP[34]/N07D[35] level of theory utilizing the Gaussian 16 software package.[36] The N07D basis set is based on the 6-31G(d) basis set and contains additional diffuse and polarization functions that have been shown to increase accuracy in the anharmonic computations of large aromatic systems such as PAHs.[37] A QFF is a truncated Taylor series expansion of the potential surface surrounding the equilibrium geometry, following the formula:
A simple Hessian cannot be constructed due to the non-linear nature of the potential, requiring non-conventional means to obtain the anharmonic vibrational spectrum. From here, two different methodologies are implemented via 2 order vibrational perturbation theory (VPT2).[38, 39, 40, 41] Using the built-in anharmonic frequency code within Gaussian 16, VPT2 is used to compute the anharmonic frequencies including up to three-quanta transitions. This allows for the inclusion of multi-mode combination bands that are imperative for an accurate analysis of the recorded experimental absorption data.
Alternatively, to properly treat the redistribution of intensity between resonant normal modes, a locally modified version of SPECTRO[42] is utilized to compute the anharmonic vibrational frequencies of transitions involving changes up to two vibrational quanta. SPECTRO has the advantage of using resonance polyads in the anharmonic computations.[43, 44] When two normal modes are close in frequency and produce a near-singularity in the VPT2 treatment, they are removed from the standard perturbation treatment and included in a resonance polyad matrix. The matrix allows for the treatment of resonance effects as well as normal modes that simultaneously participate in multiple resonances. Additionally, the polyad matrix allows for the distribution of intensities between coupled vibrations. This method has been described previously in greater detail.[25, 17, 18, 19, 21, 15, 16]
The resultant vibrational absorption stick spectrum is convoluted with a Gaussian line shape having a full-width at half-maximum (FWHM) of 0.5% of the central frequency as to match the bandwidth of FELIX.
@@ III.1 Phenylacetylene Absorption Spectrum
Figure 1 presents the experimental infrared absorption spectrum (black) of phenylacetylene in the 100-2300 cm frequency range, the 3-quanta computational anharmonic stick spectrum (purple), and the artificially broadened anharmonic spectrum (blue) based on this stick spectrum. The experimental band center, FWHM, and relative intensity along with the pertinent associated computational anharmonic mode labeling, frequency, and intensity of the pertinent bands are presented in Table 1. The full list of experimental and computational transitions are provided in Table S1 of the Supplementary Information (SI). The experimental transitions are determined via a least-squares fit using a Gaussian line shape. Overall, approximately 40 features can be identified from the experimental data above a signal-to-noise ratio of 2. Because of the resolution of the experiment, the many overlapping bands leads to unstructured, broad shapes and shoulders. Studying the computational spectrum leads to a deeper understanding as to why many of the features appear so broad and structure-less. For example, in the experimental spectrum the fit detected two neighboring features at 753.1 and 764.8 cm within the broad feature centered at 753.1 cm, which exhibits a shoulder on the high-frequency side. Using the anharmonic computations, these two features can be reliably assigned to (out-of-plane aromatic CH wag) and (in-plane aromatic ring breathing mode), respectively, clearly explaining the finer details of the 753.1 cm feature.
@@
Across the entirety of the spectrum (Figure 1), every experimental feature is assigned to either a single or multiple vibrational modes. Interestingly, the two features at 609.4 and 1217.6 cm stand out as having poor agreement with the computations. The data in Table 1 show that the feature centered at 609.4 cm is assigned to , which is the out-of-plane acetylene CH wag that causes a large change in the dipole moment and leads to high intensity. The difference between the experiment and computed position for this mode is 22.6 cm, larger than expected for a fundamental mode at this level of theory, especially considering that the surrounding features are well-described by the computations (e.g., the feature at 620.6 cm is assigned to the + combination band computed to be at 626.7 cm). Anharmonic computations of out-of-plane bending motions of aromatic systems consisting of sp hybridized carbons have been previously reported to exhibit unexpected behavior when using DFT methods in combination with some basis sets,[45]. We believe that this is indeed the cause of the issue seen here. Additionally, phenylacetylene has been reported to exhibit large couplings between various normal modes on the acetylene substituent.[31] The initial error in the potential energy surface computation for is compounded when moving to higher-order transitions along with the general decrease in accuracy when computing the frequency of overtone transitions. This leads to a larger error in the frequency for the feature centered at 1217.6 cm, which arises from the first overtone of (). The other feature that has poor agreement between experiment and computation is the feature at 2119.0 cm, assigned to (2165.3 cm), the acetylene CC stretch. It is unclear at the moment why there is such a large difference between experiment and theory for this feature.
@@
Moving to higher frequencies, each feature from 1250-2000 cm is assigned to one or more underlying modes. Due to the congested nature of this region, some features are assigned to multiple nearby modes that cannot be separated effectively. Figure 2 illustrates how important it is to include anharmonicity in the computations of the vibrational spectrum of PAHs. In the 1250-2000 cm frequency range, only seven fundamental frequencies (green) exist, making it impossible to completely assign the complicated spectrum of phenylacetylene. However, when anharmonicity is included, hundreds of relatively intense overtone and combination bands present themselves that agree quite nicely with the experimental features. Furthermore, the anharmonic intensities show better agreement with the experimental intensities than the double harmonic intensities do.
@@
Figure 3 illustrates the 2-quanta (red) and 3-quanta (blue) vibrational modes computed for phenylacetylene. The importance of including the 3-quanta modes cannot be overstated, as can be seen from the figure and is highlighted by Table 1. At least four 3-quanta modes (combination bands involving three fundamental vibrations or one fundamental mode and one overtone, as well as a second overtone mode) are assigned to features in the spectrum that would have been missed if the higher-order modes would not have been included in the computations. For example, around 1800 cm, the experiment reveals a double feature that can only be explained by including 3-quanta modes. Of note is that the 3-quanta computations are not performed with polyad matrices and therefore do not benefit from the proper treatment of intensity sharing. The less accurate intensities can be seen in the systematically larger intensity for almost all modes in the 3-quanta computations versus those of the 2-quanta in Figure 3. Not having a polyad matrix available will also affect the frequencies of the bands in the 3-quanta computations, although to a lesser degree. These issues will be addressed in subsequent studies, when the capability to compute 3-quanta modes is developed further.
@@
Figure 4 presents the absorption spectrum of d–phenylacetylene in the range of 100-2300 cm with the experiment shown in black, the anharmonic 3-quanta stick spectrum in purple, and the anharmonic stick spectrum broadened with a Gaussian lineshape (FWHM of 0.5% of the central frequency) in blue. Table 2 lists computational and experimental frequencies of prominent bands. The full list of experimental and computational frequencies is provided in Table S2 of the SI. In the experimental spectrum bands are observed throughout the spectral range with mostly sharp, well-defined peaks below 800 cm and broad, structure-less features above. The sole physical difference between phenylacetylene and d–phenylacetylene is a single deuteration, but this leads to major spectroscopic changes. Many spectral shifts originate from modes that involve motion of the deuterium atom, and adjacent modes (such as the acetylene CC stretch) are also impacted. Here, theory provides insights into what drives these spectral changes.
@@
There is generally good agreement between experiment and computation across the entire spectral range with a few notable outliers. In the range from 450-560 cm, there are four features present in the experimental spectrum while theory predicts six bands. The three largest peaks in the experimental spectrum occur at 475.7, 480.6 and 527.2 cm, with the peaks at 475.7 and 480.6 cm overlapping (Figure S1 shows an expanded view of this feature). An unambiguous assignment is difficult to make here, but the strongest peak at 480.6 cm most likely originates from , which is the in-plane acetylene CD wag. , the out-of-plane (OOP) acetylene CD wag computed at 461.5 cm, is tentatively assigned to the experimental feature at 475.7 cm. This mode is analogous to in the standard isotopologue. Deuteration of the acetylene hydrogen leads to an experimental isotopic shift of 128.8 cm for the OOP acetylene CH wag mode. Assigning the feature at 527.2 cm is more difficult due to three modes computed near the central peak position and the intensity mismatch. The carrier of this peak is probably , the OOP acetylene CC bend predicted at 528.1 cm. After the VPT2 treatment, the computed intensity of drops from 18.229 km mol (0.524 relative to ) to 4.408 km mol (0.097 relative to ). As mentioned previously, OOP bending modes are sometimes not well described by DFT/VPT2, and intensities in general are difficult to calculate. Therefore, better agreement is found when considering the double harmonic intensities. Additionally, the lack of intensity sharing in these calculations contributes to the intensity mismatch for all of the features mentioned here. The other two small peaks predicted at 535.9 (+) and 539.1 () cm add to the intensity of the central peak, and lead as well to the appearance of the shoulder on the high frequency side.
Moving to higher frequencies, there is somewhat less agreement between experiment and computation in the 900-1100 cm range. However, 10 prominent peaks are seen in both spectra, indicating the ability to understand which modes are involved without directly assigning individual peaks. There is a mixture of transitions in this region characterized by fundamental, overtone, and 2- and 3-quanta combination bands contributing to features observed here; 2 reports the largest predicted intensity (see SI Table S2). Only three fundamental transitions have appreciable intensity in this region (see SI Figure S2), indicating and reinforcing the strong impact of anharmonicity on phenylacetylene [31] and the need for anharmonic computations to properly analyze and understand its spectrum.

### 22 Feb 2024 v2

@@ Abstract
Anharmonicity strongly influences the absorption and emission spectra of polycyclic aromatic hydrocarbon (PAH) molecules. Here, IR-UV ion-dip spectroscopy experiments together with detailed anharmonic computations reveal the presence of fundamental, overtone, as well as 2- and 3-quanta combination band transitions in the far- and mid-infrared absorption spectrum of phenylacetylene and its singly deuterated isotopologue. Strong absorption features in the 400-900 cm range originate from CH(D) in-plane and out-of-plane wags and bends, as well as bending motions including the CC and CH bonds of the acetylene substituent and the aromatic ring. For phenylacetylene, every absorption feature is assigned either directly or indirectly to a single or multiple vibrational mode(s). The measured spectrum is dense, broad, and structureless in many regions but well characterized by computations. Upon deuteration, large isotopic shifts are observed. At frequencies above 1500 cm for d–phenylacetylene, a one-to-one match is seen when comparing computations and experiment with all features assigned to combination bands and overtones. The CC stretch observed in phenylacetylene is not observed in d–phenylacetylene due to a computed 40-fold drop in intensity. Overall, a careful treatment of anharmonicity that includes 2- and 3-quanta modes is found to be crucial to understand the rich details of the infrared spectrum of phenylacetylene. Based on these results, it can be expected that such an all-inclusive anharmonic treatment will also be key for unraveling the infrared spectra of PAHs in general.
@@ I Introduction
PAHs are a class of molecules composed of fused hexagonal rings of sp-hybridized carbon atoms with hydrogens decorating the periphery, which are known to be ubiquitous in space based on their infrared emission features detected throughout many interstellar regions, i.e., the aromatic infrared bands (AIBs).[1, 2] Recently, experimental and computational studies of the infrared absorption spectra of various PAHs have revealed large anharmonic effects. Previous attempts to account for anharmonicity included the use of harmonic scaling factors.[3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] However, scaling factors do not account for important higher-order transitions such as combination bands and overtones, nor do they provide a good treatment of systems with a shallow minimum in their PESs.[14]
@@
In the recent past dedicated gas-phase experiments have been performed for a number of PAHs, and these indeed reveal that for these compounds anharmonicity is essential to come to a proper interpretation of their infrared absorption spectra. For example, IR-UV ion-dip molecular beam experiments have been used to record infrared absorption spectra of neutral naphthalene, anthracene, tetracene and pentacene, showing the appearance of many combination bands in the 1600-2000 cm region.[25, 17, 18, 26, 27, 28] Similarly, molecular beam experiments on the larger PAHs coronene, peropyrene, ovalene, and hexa(peri)benzocoronene, have shown the presence of many combination bands and overtones in both the 1600–2000 cm and 2950–3150 cm regions, whereas at lower frequencies (100 and 1000 cm) fundamental transitions dominate.[29] Recently, using an infrared table-top laser system to cover the 3 µm spectral region, our research group performed ion-dip experiments on phenylacetylene and d–phenylacetylene (the singly deuterated isotopologue of phenylacetylene with deuteration taking place at the acetylene hydrogen),[16] representing an ideal model system to assess the impact of anharmonic effects on PAHs, in particular focusing on the possibility that higher-order quanta modes are a significant contributor to their infrared spectra. Remarkably, a major influence of 2- and 3-quanta modes arising from aromatic and acetylenic C-H and C-D stretching bands is observed.[30] Notably, phenylacetylene was recently detected in the dark molecular cloud TMC-1 via its pure rotational transitions,[31] making the detailed understanding of its infrared spectrum a pressing need.
@@ II.1 Experimental
Molecular beams of phenylacetylene and d–phenylacetylene are formed by placing the liquid compounds in a reservoir kept at room temperature that is seeded with a stream of Ar gas injected at a backing pressure of 4 bar. Using a Series 9 pulsed valve from General Valve, the Ar-phenylacetylene mixtures are expended into vacuum, forming a cold and directional molecular beam that is collimated by a 2 mm skimmer. The pulsed valve is operated at 10 Hz. After collimation, the molecular beam enters the ionization region of a reflectron time-of-flight mass spectrometer (R.M. Jordan D-850) equipped with a 40 mm dual microchannel plate (MCP) detector (Jordan Co. C-726) that has a mass resolution of = 2200 at 100 amu. In previous studies on the same apparatus with similarly-sized molecular systems, molecular dynamic simulations were used to compute infrared spectra at different temperatures, finding good agreement with a vibrational temperature of 50 K.[14] Similarly, previous estimations of the molecular beam temperature place its value in the range from 10 to 50 K.[32, 33] Values much higher than this range would lead to visible hot bands in REMPI scans, which is not the case in the spectra in this study.
@@ II.2 Computational
Computation of the optimized geometry; harmonic normal modes; and quadratic, cubic, and quartic normal coordinate force constants (quartic force field; QFF) is performed at the B3LYP[35]/N07D[36] level of theory utilizing the Gaussian 16 software package.[37] The computations were done using the verytight optimization criteria and a custom integration grid consisting of 200 radial shells and 974 angular points per shell (compared to 99 radial shells and 590 angular points per shell in the default UltraFine grid) The N07D basis set is based on the 6-31G(d) basis set and contains additional diffuse and polarization functions that have been shown to increase accuracy in the anharmonic computations of large aromatic systems such as PAHs.[38] A QFF is a truncated Taylor series expansion of the potential surface surrounding the equilibrium geometry, following the formula:
The QFF is computed in normal mode coordinates via a linear relationship to produce a Cartesian coordinate QFF. Additionally, semi-diagonal quartic terms are employed as is the default in Gaussian16.[39]
From here, two different methodologies are implemented via 2 order vibrational perturbation theory (VPT2).[40, 41, 42, 43] Using the built-in anharmonic frequency code within Gaussian 16, VPT2 is used to compute the anharmonic frequencies including up to three-quanta transitions. This allows for the inclusion of multi-mode combination bands that are imperative for an accurate analysis of the recorded experimental absorption data.
Alternatively, to properly treat the redistribution of intensity as a result of accidental resonances between vibrational states, a locally modified version of SPECTRO[44] is utilized to compute the anharmonic vibrational frequencies of transitions involving changes up to two vibrational quanta. SPECTRO has the advantage of using resonance polyads in the anharmonic computations.[45, 46] When two vibrational states are close in frequency and produce a near-singularity in the VPT2 treatment, they are removed from the standard perturbation treatment and included in a resonance polyad matrix. The matrix allows for the treatment of resonance effects as well as normal modes that simultaneously participate in multiple resonances. Additionally, the polyad matrix allows for the distribution of intensities between coupled vibrational states. Vibrational modes with frequencies below 300 cm are excluded from the SPECTRO VPT2 calculation. This method has been described previously in greater detail.[25, 17, 18, 19, 21, 15, 16]
The resultant vibrational absorption stick spectrum is convoluted with a Gaussian line shape having a full-width at half-maximum (FWHM) of 0.5% of the computational frequency as to match the bandwidth of FELIX.
@@ III.1 Phenylacetylene Absorption Spectrum
Figure 1 presents the experimental infrared absorption spectrum (black) of phenylacetylene in the 100-2300 cm frequency range, the harmonic stick spectrum (green), the 3-quanta anharmonic stick spectrum (purple), and the artificially broadened anharmonic spectrum (blue) based on this stick spectrum. The experimental band center, FWHM, and relative intensity along with the pertinent associated computational anharmonic mode labeling, frequency, and intensity of the pertinent bands are presented in Table 1. The full list of computational frequencies and experimental transition assignments are provided in Table S1 of the Supplementary Information (SI). The experimental transitions are determined via a least-squares fit using a Gaussian line shape. Overall, approximately 40 features can be identified from the experimental data above a signal-to-noise ratio of 2. Because of the resolution of the experiment, the many overlapping bands leads to unstructured, broad shapes and shoulders. Studying the computational spectrum leads to a deeper understanding as to why many of the features appear so broad and structure-less. For example, in the experimental spectrum the fit detected two neighboring features at 753.1 and 764.8 cm within the broad feature centered at 753.1 cm, which exhibits a shoulder on the high-frequency side. Using the anharmonic computations, these two features can be reliably assigned to (out-of-plane aromatic CH wag) and (in-plane aromatic ring breathing mode), respectively, clearly explaining the finer details of the 753.1 cm feature.
@@
Across the entirety of the spectrum (Figure 1), every experimental feature is assigned to either a single or multiple vibrational modes. Interestingly, the two features at 609.4 and 1217.6 cm stand out as having poor agreement with the computations. The data in Table 1 show that the feature centered at 609.4 cm is assigned to , which is the out-of-plane (OOP) acetylene CH wag that causes a large change in the dipole moment and leads to high intensity. The difference between the experiment and computed position for this mode is 22.6 cm, larger than expected for a fundamental mode at this level of theory, especially considering that the surrounding features are well-described by the computations (e.g., the feature at 620.6 cm is assigned to the + combination band computed to be at 626.7 cm). Anharmonic computations of OOP bending motions of aromatic systems consisting of sp hybridized carbons have been previously reported to exhibit unexpected behavior when using DFT methods in combination with some basis sets,[47]. In this study, different DFT and correlated methods with varied basis sets were tested and all suffered the same issues with the OOP bending mode, and in fact, the B3LYP/N07D method performed the best against experiment. Therefore, we believe that this is indeed the cause of the issue seen here. To confirm that there are no issues with basis set size, the harmonic vibrational spectrum of phenylacetylene calculated with B3LYP/N07D compares well to those calculated with B3LYP/6-311++G(d,p), agreeing to 16 cm in the CH stretch region and approximately 4 cm for lower frequency modes. The in-plane acetylene CH wag is slightly more sensitive to basis set size; however, a difference of approximately 3% will not impact the results in this study. Additionally, phenylacetylene has been reported to exhibit large couplings between various vibrational states of the acetylene substituent.[30] The initial error in the potential energy surface computation for is compounded when moving to higher-order transitions along with the general decrease in accuracy when computing the frequency of overtone transitions (see Ref. [48] for more details). This leads to a larger error in the frequency for the feature centered at 1217.6 cm, which arises from the first overtone of (). The other feature that has poor agreement between experiment and computation is the feature at 2119.0 cm, assigned to (2165.3 cm), the acetylene CC stretch. It is unclear at the moment why there is such a large difference between experiment and theory for this feature.
@@
Moving to higher frequencies, each feature from 1250-2000 cm is assigned to one or more underlying modes. Due to the congested nature of this region, some features are assigned to multiple nearby modes that cannot be separated effectively. The harmonic stick spectrum (green) in Figure 1 illustrates how important it is to include anharmonicity in the computations of the vibrational spectrum of PAHs. In the 1250-2000 cm frequency range, only seven fundamental frequencies (green) exist, making it impossible to completely assign the complicated spectrum of phenylacetylene. However, when anharmonicity is included, hundreds of relatively intense overtone and combination bands present themselves that agree quite nicely with the experimental features. Furthermore, the anharmonic intensities show better agreement with the experimental intensities than the double harmonic intensities do.
@@
Figure 2 illustrates the 2-quanta (red) and 3-quanta (blue) vibrational modes computed for phenylacetylene. The importance of including the 3-quanta modes cannot be overstated, as can be seen from the figure and is highlighted by Table 1. At least four 3-quanta modes (combination bands involving three fundamental vibrations or one fundamental mode and one overtone, as well as a second overtone mode) are assigned to features in the spectrum that would have been missed if the higher-order modes would not have been included in the computations. For example, around 1800 cm, the experiment reveals a double feature that can only be explained by including 3-quanta modes. Of note is that the 3-quanta computations are not performed with polyad matrices and therefore do not benefit from the proper treatment of intensity sharing. The less accurate intensities can be seen in the systematically larger intensity for almost all modes in the 3-quanta computations versus those of the 2-quanta in Figure 2. A glaring example of this is the disappearance of the 2 feature in the 3-quanta computation. As the result of many strong resonance coupling interactions, the intrinsic intensity of this mode is redistributed over various nearby features. Not utilizing a polyad matrix will also affect the frequencies of the bands in the 3-quanta computations, although to a lesser degree. These issues will be addressed in subsequent studies, when the capability to compute 3-quanta modes is developed further.
@@
Figure 3 presents the absorption spectrum of d–phenylacetylene in the range of 100-2300 cm with the experiment shown in black, the harmonic stick spectrum in green, the anharmonic 3-quanta stick spectrum in purple, and the anharmonic stick spectrum broadened with a Gaussian lineshape (FWHM of 0.5% of the computational frequency) in blue. Table 2 lists computational and experimental frequencies of prominent bands. The full list of experimental and computational frequencies is provided in Table S2 of the SI. In the experimental spectrum bands are observed throughout the spectral range with mostly sharp, well-defined peaks below 800 cm and broad, structure-less features above. The sole physical difference between phenylacetylene and d–phenylacetylene is a single deuteration, but this leads to major spectroscopic changes. Many spectral shifts originate from modes that involve motion of the deuterium atom, and adjacent modes (such as the acetylene CC stretch) are also impacted. Here, theory provides insights into what drives these spectral changes.
@@
There is generally good agreement between experiment and computation across the entire spectral range with a few notable outliers. In the range from 450-560 cm, there are four features present in the experimental spectrum while theory predicts six bands. The three largest peaks in the experimental spectrum occur at 475.7, 480.6 and 527.2 cm, with the peaks at 475.7 and 480.6 cm overlapping (Figure S1 shows an expanded view of this feature). An unambiguous assignment is difficult to make here, but the strongest peak at 480.6 cm most likely originates from , which is the in-plane acetylene CD wag. , the OOP acetylene CD wag computed at 461.5 cm, is tentatively assigned to the experimental feature at 475.7 cm. This mode is analogous to in the standard isotopologue. Deuteration of the acetylene hydrogen leads to an experimental isotopic shift of 128.8 cm for the OOP acetylene CH wag mode. Assigning the feature at 527.2 cm is more difficult due to three modes computed near the central peak position and the intensity mismatch. The carrier of this peak is probably , the OOP acetylene CC bend predicted at 528.1 cm. After the VPT2 treatment, the computed intensity of drops from 18.229 km mol (0.524 relative to ) to 4.408 km mol (0.097 relative to ). As mentioned previously, OOP bending modes are sometimes not well described by DFT/VPT2, and intensities in general are difficult to calculate. Therefore, better agreement is found when considering the double harmonic intensities. Additionally, the lack of intensity sharing in these calculations contributes to the intensity mismatch for all of the features mentioned here. The other two small peaks predicted at 535.9 (+) and 539.1 () cm add to the intensity of the central peak, and lead as well to the appearance of the shoulder on the high frequency side.
Moving to higher frequencies, there is somewhat less agreement between experiment and computation in the 900-1100 cm range. However, 10 prominent peaks are seen in both spectra, providing the ability to understand which modes are involved without directly assigning individual peaks. There is a mixture of transitions in this region characterized by fundamental, overtone, and 2- and 3-quanta combination bands contributing to features observed here; 2 reports the largest predicted intensity (see SI Table S2). Only three fundamental transitions have appreciable intensity in this region (harmonic sticks in Figure 3), indicating and reinforcing the strong impact of anharmonicity on phenylacetylene [30] and the need for anharmonic computations to properly analyze and understand its spectrum.